Plz help me to do both quetions.6. The continuous random variables X and Y have joint density 2e2)for and y 20 otherwise Find P(X >Y). (Answer: 3) 7. The continuous random variables X and Y have joint probability density function 10y for 01 and 0 yr fzy(z,y)=ï o otherwise Find the marginal density function fy(y). Show your work. Do not forget to indicate where the density is non-zero.
Answer #1

6)

P(X>Y)=\int_{0}^{\infty }\int_{0}^{x } f(x,y) dy dx =\int_{0}^{\infty }\int_{0}^{x } 2e-xe-2y dy dx =\int_{0}^{\infty } 2e-x(-e-2y/2)|x0 dx

=\int_{0}^{\infty } e-x(1-e-2x) dx =\int_{0}^{\infty }(e-x -e-3x) dx =(-e-x+e-3x/3)|\infty0 =(1-1/3) =2/3

7)

fy(y) =\int_{y}^{1} f(x,y) dx =\int_{y}^{1} 10x2y dx = 10x3y/3 |1y =(10/3)*y*(1-y3) for 0<=y<=1